The electric cooperative needs to know the mean household usage of electricity by its non-commercial customers in kWh per day. They would like the estimate to have a maximum error of 0.14 kWh. A previous study found that for an average family the standard deviation is 2.4 kWh and the mean is 16.3 kWh per day. If they are using a 90% level of confidence, how large of a sample is required to estimate the mean usage of electricity? Round your answer up to the next integer.
Answer #1

90% level of confidence = 1.645

Confidence Level Area between 0 and z-score Z-score 80% 0.4000 1.282 90% 0.4500 1.645 95% 0.4750 1.960 98% 0.4900 2.326

Now, find the margin of error M as such

M = z * (σ / sqrt(n))

0.14 = 1.645 * (2.4/ sqrt(n))

0.14*sqrt(n) = 1.645*2.4

sqrt(n) = (1.645*2.4) / 0.14

n = ((1.645*2.4) / 0.14)^2

n = 795.24

how large of a sample is required to estimate the mean usage of electricity?

795